Solve the following equation for x:
5(x – 2)^2 = 3
Express your answer(s) in both "simplified radical" form, and in decimal form.
I came up with x=2+or-sqrt(3/5). Could someone please check this one for me. Thanks.

Answer by Nate(3500)

5(x – 2)^2 = 3
(x – 2)^2 = 3/5
x - 2 = +- sqrt(3/5)
x = 2 +- sqrt(3/5)
x = 2 +- sqrt(3)/sqrt(5)
x = 10/5 +- sqrt(15)/5
x = (10 +- sqrt(15))/5
x = 2.77 or x = 1.23